# Towards Thinking-Optimal Scaling of Test-Time Compute for LLM Reasoning

### v1

@@ Abstract
Recent studies have shown that making a model spend more time thinking through longer Chain of Thoughts (CoTs) enables it to gain significant improvements in complex reasoning tasks. While current researches continue to explore the benefits of increasing test-time compute by extending the CoT lengths of Large Language Models (LLMs), we are concerned about a potential issue hidden behind the current pursuit of test-time scaling: Would excessively scaling the CoT length actually bring adverse effects to a model’s reasoning performance? Our explorations on mathematical reasoning tasks reveal an unexpected finding that scaling with longer CoTs can indeed impair the reasoning performance of LLMs in certain domains. Moreover, we discover that there exists an optimal scaled length distribution that differs across different domains. Based on these insights, we propose a Thinking-Optimal Scaling strategy. Our method first uses a small set of seed data with varying response length distributions to teach the model to adopt different reasoning efforts for deep thinking. Then, the model selects its shortest correct response under different reasoning efforts on additional problems for self-improvement. Our self-improved models built upon Qwen2.5-32B-Instruct outperform other distillation-based 32B o1-like models across various math benchmarks, and achieve performance on par with QwQ-32B-Preview.
@@ 1 Introduction
Behind the promising paradigm of test-time scaling, there is a few concurrent studies [3, 20] highlighting an efficiency issue of overthinking in existing o1-like models, where they tend to generate an excessive number of tokens, even for simple questions that could be answered correctly with just a few tokens. However, we are concerned about a more critical issue that could the excessive pursuit of longer CoTs have negative impacts on the model’s reasoning performance? That is, besides the efficiency issues, our aim is to explore and study whether and how overly test-time scaling could potentially impair the reasoning performance of LLMs, typically in the math domain.
To study the problem, we first calculate and compare the accuracies and used reasoning tokens of several o1-like models and their corresponding System-1 thinking models on MATH500 [19] and AIME2024111https://huggingface.co/datasets/AI-MO/aimo-validation-aime (see Figure 1). We find that subsequent o1-like models, QwQ-32B-Preview [27] as an typical example, generate much more tokens but gain only limited improvements in model performance. This preliminary exploration indicates that scaling to more reasoning tokens might not consistently lead to better performance. Then, to fairly investigate the performance comparison of the same base model after scaling with different lengths of CoTs, we conduct additional experiments on LLaMA3.1-8B-Instruct [21] and Qwen2.5-32B-Instruct [26]. Specifically, we utilize QwQ-32B-Preview [27] to generate and filter three types of reasoning paths with different lengths for the same set of prompts. Then, we teach the base model to use different reasoning efforts (i.e., different numbers of reasoning tokens) to solve a given problem based on learning on different subsets. Surprisingly, we find that training with longer reasoning paths leads to worse performance especially in easier tasks, and there exists an optimal reasoning effort that varies across tasks of different difficulty levels. Our further analysis reveals that longer CoTs may contain more erroneous steps. Though including a certain incorrect steps and subsequent reflective steps can teach the model how to correct errors during inference, excessive erroneous steps can have a negative impact.
@@ 2 Related Work
LLM Reasoning Leveraging the Chain-of-Thought (CoT) technique [36, 44], LLMs have demonstrated impressive performance on various reasoning tasks [30, 29, 31]. CoT enables LLMs to decompose the entire problem into several sub-goals and then reason step-by-step to achieve a more reliable answer. Among various LLM reasoning tasks, mathematical reasoning has become one of the most widely studied and important tasks. Current work on LLM math reasoning primarily focuses on: synthesizing large-scale and diverse math data [42, 41, 17], constructing challenging math reasoning benchmarks [8, 9], training powerful process reward models (PRMs) [19, 34, 32], and designing more effective alignment or reinforcement learning algorithms to improve math reasoning capabilities of LLMs [16, 12, 5].
Test-Time Scaling Recently, scaling test-time compute of LLMs has shown significant potential for further improving their reasoning performance [1, 2]. Existing test-time scaling studies can be divided into several categories: (1) Sampling-based scaling aims to increase the number of individual reasoning paths of LLMs when solving a given problem. Then the most reliable answer is selected from all the generated options using mechanisms such as majority voting [35], weighted majority voting [18], or best-of-N selection [19]. (2) Tree search-based scaling expands reasoning paths by constructing tree-like trajectories, allowing LLMs to explore diverse options at each state and continue reasoning along the most promising directions. Tree-of-Thoughts (ToT) [40] and Monte Carlo Tree Search (MCTS) [37, 43, 33, 45] are two typical tree search-based test-time scaling methods. (3) In-context search-based scaling enables LLMs to learn to search, backtrack and re-explore within one single CoT path [7]. Recently, OpenAI’s o1 model [25] have made a significant breakthrough in this line. It leverages reinforcement learning to scale the lengths of CoT to enable LLMs to perform thorough thinking through reflection, verification and re-exploration when solving problems. Following the same line, a series of studies [32, 6, 27, 11, 22, 14] have been proposed to scale CoT lengths during inference time. Our work also primarily focuses on the scaling properties of o1-like models.
@@
We notice that there is a few concurrent studies [3, 20] highlighting that existing o1-like models exhibit overthinking issues, often generating an excessive number of tokens for simple problems with minimal benefit. Thus, they aim to shortening the CoT lengths of o1-like models while preserving their performance. However, our work differs in that we aim to uncover a deeper and more critical issue: scaling with more tokens can, in some cases, even degrade the model’s performance. Thus, our work focuses on achieving optimal test-time scaling from base models in both aspects of effectiveness and efficiency.
3.1 Preliminary Analysis on Existing o1-Like Models
Though o1-like models has proven to be much more effective on reasoning tasks than previous System-1 thinking models, we are curious about the scaling process behind the these o1-like models. That is, we want to explore that: How effective has their scaling achieved compared to their corresponding System-1 thinking models(e.g., QwQ-32B-Instruct v.s. Qwen2.5-32B-Instruct)?
We first conduct a preliminary analysis on several existing typical o1-like models along with their corresponding System-1 thinking models. Specifically, we choose o1-mini [25] v.s. gpt-4o/4o-mini [24], [^2] Note that o1-mini and 4o/4o-mini do not have equivalent number of parameters, but we make a rough comparison here. Gemini2.0-Flash-Thinking-Exp.-1219 [11] v.s. Gemini2.0-Flash-Exp. [10], and QwQ-32B-Preview [27] v.s. Qwen2.5-32B-Instruct [26] as our experimental models. We calculate the accuracy and the average number of generated tokens of each model on two typical benchmarks: MATH500 [19]: 500 high school math competition problems across various subjects, sampled from MATH benchmark [13]; AIME2024: 30 challenging problems from the American Invitational Mathematics Examination (AIME). To address the issue of token counts not being directly comparable due to the different tokenizers used by different models, we standardize by using Qwen2.5 tokenizer to tokenize the reasoning completions of different models and then calculate the number of tokens. As the internal CoT of o1-mini is not available to users, we use an estimation strategy based on the summary part, the number of reasoning tokens and total number of completion tokens returned from the o1-mini model to estimate the number of tokens of hidden CoT tokenized by Qwen2.5 tokenizer. Details are in Appendix A. We set the maximum number of generation tokens to 16,384 for each model in all evaluations in this paper.
@@ 3.2 Deeper Explorations on The Scaling Process of CoT Length
Model #Problems #Tokens
@@
The above analysis still faces a problem that the base models of different o1-like models are not identical, making it unfairly to compare the impacts of scaled CoT lengths on test-time scaling effects of different models. Therefore, we conduct experiments on LLaMA3.1-8B-Instruct [21] and Qwen2.5-32B-Instruct [26] to fairly investigate the impact of the scaling efforts on the effectiveness of test-time scaling. Specifically, we first use three system prompts (refer to Figure 2), corresponding to different levels of reasoning effort (“Low”, “Medium” and “High”), to prompt QwQ-32B-Preview to generate solutions of different numbers of tokens for the same set of math problems sampled from NuminaMath [17]. We then filter out the problems that can be answered correctly under all three reasoning efforts, along with the corresponding three reasoning paths of different lengths. However, we find that QwQ-32B-Preview has relatively poor instruction-following abilities, reflected in that the length distributions of the generated responses for the same question does not closely match the specified system prompts. Therefore, for a given problem, we further reorder the three responses based on their lengths and keep them if their pairwise length difference consistently exceeds 300 tokens. The length is determined either by LLaMA3.1 tokenizer or Qwen2.5 tokenizer depending on the chosen experimental model. Finally, we curate a set of 1.3K problems, each accompanied by three o1-like responses of varying lengths. The data statistics of each set for each model is shown in Table 1. We assign different system prompts (the same in Figure 2) to each type of responses and train the base model on all three types of samples. By doing so, we ensure the consistency between the base model and the training problems, allowing for a fair comparison on the impacts of different scaling efforts on the effectiveness of test-time scaling. We refer to the fine-tuned model as the “tag" model (LLaMA3.1-8B-Tag and Qwen2.5-32B-Tag). During inference, we can use different system prompts to guide the tag model in applying varying levels of reasoning effort to answer the questions.
For training, the learning rate is $1\times 10^{-5}$, the batch size is $32$, the number of epochs is 3 for Qwen2.5-32B-Tag and 5 for LLaMA3.1-8B-Tag. For evaluation, besides MATH500 and AIME2024 introduced before, we further include GSM8K [4] that contains 1319 grade school math word problems. In the following, unless otherwise specified, we set the decoding temperature to 1.0 for o1-like models, by following the recommended setting for o1 model, [^3] https://platform.openai.com/docs/guides/reasoning which we also find is more suitable for o1-like model’s reasoning. For System-1 thinking models LLaMA3.1-8B-Instruct and Qwen2.5-32B-Instruct, we report the results for both decoding temperatures, 1.0 and 0.0, for comprehensive comparison. Each result under 1.0 temperature is averaged over 5 random seeds. The full results on LLaMA3.1-8B-Instruct and Qwen2.5-32B-Instruct are shown in Figure 3 and Figure 4 respectively. We can draw several interesting conclusions from these results: (1) A small number o1-like responses is already highly effective in enhancing the reasoning performance of LLMs. This is also consistent with the findings in previous studies [22, 14]. (2) Scaling with longer CoTs can bring negative effects to the model’s reasoning performance in certain domains, especially on easy tasks. For example, both LLaMA3.1-8B-Tag and Qwen2.5-32B-Tag perform worse under high reasoning effort compared to the other two reasoning efforts, while consuming significantly more tokens, particularly on GSM8K and MATH500. (3) There exists an optimal reasoning efforts that varies across different tasks of varying difficulty levels. As we can see, low reasoning effort consistently works best on GSM8K, while medium and high reasoning efforts are more beneficial for harder question.
@@ 3.3 Analysis on The Adverse Effects of Excessive Length Scaling
Here, we take a deeper step to explore why training on longer CoTs leads to a decline in model’s reasoning performance. We randomly selected 100 problems from the tag model’s training set along with their responses under three different reasoning efforts. We first follow the existing study [3] to use gpt-4o to determine the total number of reasoning rounds contained in each response. Each reasoning round is defined as a complete reasoning process or verification process that contains a final answer. Besides, we further use gpt-4o to determine the number of reasoning rounds that contain erroneous steps or wrong final answers. The utilized prompt is shown in Appendix B. We visualize the average number of reasoning rounds and the average number of erroneous reasoning rounds on each problem under each reasoning effort in Figure 5.
First, we can see that the number of reasoning rounds consistently increases from low reasoning effort to high reasoning effort. It could lead to the overthinking issue in reasoning models [3], where redundant tokens are used to answer the question through repetitive verification. This, however, should not affect the accuracy if all reasoning rounds are correct. Unfortunately, we observe a pattern that the number of erroneous reasoning rounds also increases when the reasoning effort becomes higher. Training the model on more wrong steps would bring adverse effects to the model’s reasoning abilities, which can explain why scaling with high reasoning effort leads to worse results. Therefore, we can conclude that while including a certain incorrect and reflective steps can help the model learn to correct errors during inference, an excess of erroneous steps can have a detrimental impact on model’s learning.

### v2

@@ Abstract
Recent studies have shown that making a model spend more time thinking through longer Chain of Thoughts (CoTs) enables it to gain significant improvements in complex reasoning tasks. While current researches continue to explore the benefits of increasing test-time compute by extending the CoT lengths of Large Language Models (LLMs), we are concerned about a potential issue hidden behind the current pursuit of test-time scaling: Would excessively scaling the CoT length actually bring adverse effects to a model’s reasoning performance? Our explorations on mathematical reasoning tasks reveal an unexpected finding that scaling with longer CoTs can indeed impair the reasoning performance of LLMs in certain domains. Moreover, we discover that there exists an optimal scaled length distribution that differs across different domains. Based on these insights, we propose a Thinking-Optimal Scaling strategy. Our method first uses a small set of seed data with varying response length distributions to teach the model to adopt different reasoning efforts for deep thinking. Then, the model selects its shortest correct response under different reasoning efforts on additional problems for self-improvement. Our self-improved models built upon Qwen2.5-32B-Instruct outperform other distillation-based 32B o1-like models across various math benchmarks, and achieve performance on par with the teacher model QwQ-32B-Preview that produces the seed data. [^1] Code, data and models are available at https://github.com/RUCBM/TOPS.
@@ 1 Introduction
Behind the promising paradigm of test-time scaling, there is a few concurrent studies [3, 22] highlighting an efficiency issue of overthinking in existing o1-like models, where they tend to generate an excessive number of tokens, even for simple questions that could be answered correctly with just a few tokens. However, we are concerned about a more critical issue that could the excessive pursuit of longer CoTs have negative impacts on the model’s reasoning performance? That is, besides the efficiency issues, we aim to explore and study whether and how overly test-time scaling could potentially impair the reasoning performance of LLMs, typically in the math domain.
To study the problem, we first calculate and compare the accuracies and used reasoning tokens of several o1-like models and their corresponding System-1 thinking models on MATH500 [20] and AIME2024222https://huggingface.co/datasets/AI-MO/aimo-validation-aime (see Figure 1). We find that subsequent o1-like models, QwQ-32B-Preview [30] as an typical example, generate much more tokens but gain only limited improvements in model performance. This preliminary exploration indicates that scaling to more reasoning tokens might not consistently lead to better performance. Then, to fairly investigate the performance comparison of the same base model after scaling with different lengths of CoTs, we conduct additional experiments on LLaMA3.1-8B-Instruct [24] and Qwen2.5-32B-Instruct [29]. Specifically, we utilize QwQ-32B-Preview [30] to generate and filter three types of reasoning paths with different lengths for the same set of prompts. Then, we teach the base model to use different reasoning efforts (i.e., different numbers of reasoning tokens) to solve a given problem based on learning on different subsets. Surprisingly, we find that training with longer reasoning paths leads to worse performance especially in easier tasks, and there exists an optimal reasoning effort that varies across tasks of different difficulty levels. Our further analysis reveals that longer CoTs may contain more erroneous steps. Though including a certain incorrect steps and subsequent reflective steps can teach the model how to correct errors in inference, training on excessive erroneous steps can have a negative impact.
@@ 2 Related work
LLM Reasoning Leveraging the Chain-of-Thought (CoT) technique [42, 51], LLMs have demonstrated impressive performance on various reasoning tasks [34, 33, 35]. CoT enables LLMs to decompose the entire problem into several sub-goals and then reason step-by-step to achieve a more reliable answer. Among various LLM reasoning tasks, mathematical reasoning has become one of the most widely studied and important tasks. Current work on LLM math reasoning primarily focuses on: synthesizing large-scale and diverse math data [49, 48, 17], constructing challenging math reasoning benchmarks [8, 9], training powerful process reward models (PRMs) [20, 39, 36], and designing more effective algorithms to improve math reasoning capabilities of LLMs [16, 12, 5].
Test-Time Scaling Recently, scaling test-time compute of LLMs has shown significant potential for further improving their reasoning performance [1, 2]. Existing test-time scaling studies can be divided into several categories: (1) Sampling-based scaling aims to increase the number of individual reasoning paths of LLMs when solving a given problem. Then the most reliable answer is selected from all the generated options using mechanisms such as majority voting [40], weighted majority voting [18], or best-of-N selection [20]. (2) Tree search-based scaling expands reasoning paths by constructing tree-like trajectories, allowing LLMs to explore diverse options at each state and continue reasoning along the most promising directions. Tree-of-Thoughts (ToT) [47] and Monte Carlo Tree Search (MCTS) [43, 50, 37, 52] are two typical tree search-based test-time scaling methods. (3) In-context search-based scaling enables LLMs to learn to search, backtrack and re-explore within one single CoT path [7]. Recently, OpenAI’s o1 model [28] have made a significant breakthrough in this line. It leverages reinforcement learning to scale the lengths of CoT to enable LLMs to perform thorough thinking through reflection, verification and re-exploration when solving problems. Following the same line, a series of studies [36, 6, 30, 11, 25, 14, 19, 46] have been proposed to scale CoT lengths during inference time. Our work also primarily focuses on the scaling properties of o1-like models.
We notice that there is a few concurrent studies [3, 22] highlighting that existing o1-like models exhibit overthinking issues, often generating an excessive number of tokens for simple problems with minimal benefit. Thus, they aim to shorten the CoT lengths of o1-like models while preserving their performance. However, our work differs in that we aim to uncover a deeper and more critical issue: scaling with more tokens can, in some cases, even degrade the model’s performance. Thus, our work focuses on achieving optimal test-time scaling from base models in both aspects of effectiveness and efficiency.
3.1 Preliminary analysis on existing o1-like models
@@
Though o1-like models has proven to be much more effective on reasoning tasks than previous System-1 thinking models, we are curious about the scaling process behind the these o1-like models. That is, we want to explore that: How effective has their scaling achieved compared to their corresponding System-1 thinking models(e.g., QwQ-32B-Preview v.s. Qwen2.5-32B-Instruct)?
We first conduct a preliminary analysis on several existing typical o1-like models along with their corresponding System-1 thinking models. Specifically, we choose o1-mini [28] v.s. gpt-4o/4o-mini [27], [^3] Note that o1-mini and 4o/4o-mini do not have equivalent number of parameters, but we make a rough comparison here. Gemini2.0-Flash-Thinking-Exp.-1219 [11] v.s. Gemini2.0-Flash-Exp. [10], QwQ-32B-Preview [30] v.s. Qwen2.5-32B-Instruct [29], and DeepSeek-R1 [6] v.s. DeepSeek-V3 [21] as our experimental models. We calculate the accuracy and the average number of generated tokens of each model on two typical benchmarks: MATH500 [20]: 500 high school math competition problems across various subjects, sampled from MATH benchmark [13]; AIME2024: 30 challenging problems from the American Invitational Mathematics Examination (AIME). To address the issue of token counts not being directly comparable due to the different tokenizers used by different models, we standardize by using Qwen2.5 tokenizer to tokenize the reasoning completions of different models and then calculate the number of tokens. As the internal CoT of o1-mini is not available to users, we use an estimation strategy based on the summary part, the number of reasoning tokens and total number of completion tokens returned from the o1-mini model to estimate the number of tokens of hidden CoT tokenized by Qwen2.5 tokenizer. Details and further discussions are in Appendix B. We set the maximum number of generation tokens to 16,384 for each model in all evaluations.
@@ 3.2 Deeper explorations on the scaling process of CoT length
Model
#Problems
#Tokens
@@
The above analysis still faces a problem that the base models of different o1-like models are not identical, making it unfairly to compare the impacts of scaled CoT lengths on test-time scaling effects of different models. Therefore, we conduct experiments on LLaMA3.1-8B-Instruct [24] and Qwen2.5-32B-Instruct [29] to fairly investigate the impact of the scaling efforts on the effectiveness of test-time scaling. Specifically, we first use three system prompts (refer to Figure 7), corresponding to different levels of reasoning effort (“Low”, “Medium” and “High”), to prompt QwQ-32B-Preview to generate solutions of different numbers of tokens for the same set of math problems sampled from NuminaMath [17]. We then filter out the problems that can be answered correctly under all three reasoning efforts, along with the corresponding three reasoning paths of different lengths. However, we find that QwQ-32B-Preview has relatively poor instruction-following abilities, reflected in that the length distributions of the generated responses for the same question does not closely match the specified system prompts (refer to the empirical evidence in Appendix G). Therefore, for a given problem, we further reorder the three responses based on their lengths and keep them if their pairwise length difference consistently exceeds 300 tokens. The length is determined either by LLaMA3.1 tokenizer or Qwen2.5 tokenizer depending on the chosen experimental model. Finally, we curate a set of 1.3K problems, each accompanied by three o1-like responses of varying lengths. The data statistics of each set for each model is shown in Table 1. We assign different system prompts (the same in Figure 7) to each type of responses and train the base model on all three types of samples. By doing so, we ensure the consistency between the base model and the training problems, allowing for a fair comparison on the impacts of different scaling efforts on the effectiveness of test-time scaling. We refer to the fine-tuned model as the “tag” model (LLaMA3.1-8B-Tag and Qwen2.5-32B-Tag). During inference, we can use different system prompts to guide the tag model in applying varying levels of reasoning effort to answer the questions.
The detailed training settings are put in Appendix E.1. For evaluation, besides MATH500 and AIME2024 introduced before, we further include GSM8K [4] that contains 1319 grade school math word problems. In the following, unless otherwise specified, we set the decoding temperature to 1.0 for all o1-like models, by following the recommended setting for o1 model. [^4] https://platform.openai.com/docs/guides/reasoning For System-1 thinking models LLaMA3.1-8B-Instruct and Qwen2.5-32B-Instruct, we report the results for both decoding temperatures with 1.0 and 0.0 for comprehensive comparison. Each result under 1.0 temperature is averaged over 5 random seeds. The full results on LLaMA3.1-8B-Instruct and Qwen2.5-32B-Instruct are shown in Figure 2 and Figure 3 respectively. We also display the performance of QwQ-32B-Preview on all evaluation benchmarks when directly prompted with different prompts in Appendix H for reference. We can draw several interesting conclusions from these results: (1) A small number o1-like responses is already highly effective in enhancing the reasoning performance of LLMs. This is also consistent with the previous findings [25, 14]. (2) Scaling with longer CoTs can bring negative effects to the model’s reasoning performance in certain domains, especially on easy tasks. For example, both LLaMA3.1-8B-Tag and Qwen2.5-32B-Tag perform worse under high reasoning effort compared to the other two reasoning efforts, while consuming significantly more tokens, particularly on GSM8K and MATH500. (3) There exists an optimal reasoning efforts that varies across different tasks of varying difficulty levels. As we can see, low reasoning effort consistently works best on GSM8K, while medium and high reasoning efforts are more beneficial for harder question. Furthermore, we display the breakdown results of each model on MATH500 categorized by different problem levels in Appendix F, and the results also show the consistent conclusions on the negative effect of excessively scaling CoT lengths. We also conduct additional experiments on general reasoning tasks beyond mathematical reasoning to demonstrate that the above findings hold true in other tasks as well. The detailed results and discussions are in Appendix J.
@@
Table 2: The distribution of distinct answers under different reasoning efforts along with the average accuracy. We highlight the best accuracy and the lowest average number of distinct answers per prompt.
Model GSM8K MATH500 AIME2024
Accuracy #Answers Accuracy #Answers Accuracy #Answers
   LLaMA3.1 models
LLaMA3.1-8B-Tag-Low 87.26 1.37 59.00 2.34 07.33 4.27
LLaMA3.1-8B-Tag-Medium 87.06 1.42 61.12 2.54 07.33 4.27
LLaMA3.1-8B-Tag-High 86.89 1.47 59.36 2.59 10.00 4.00
   Qwen2.5 models
Qwen2.5-32B-Tag-Low 95.53 1.33 90.60 1.39 34.67 3.20
Qwen2.5-32B-Tag-Medium 94.33 1.14 91.48 1.36 42.00 3.13
Qwen2.5-32B-Tag-High 93.31 1.15 90.56 1.40 41.33 3.30
To further investigate the impact of scaling with varying CoT length distributions, we calculate the distributions of answers of reasoning effort-conditioned models under different reasoning efforts. Specifically, we calculate the average number of distinct answers in 5 samples per prompt under each reasoning effort. The results are in Table 2. We also display the accuracy on each benchmark for reference. We find a very interesting phenomenon. Generally, the reasoning effort that achieves the best performance on a certain benchmark also leads to the lowest average number of distinct answers per prompt. This indicates that, under the optimal thinking effort, the model can generate the most consistent answers across multiple samplings without either underthinking or overthinking.
@@ 3.3 Analysis on the adverse effects of excessive length scaling
Here, we take a deeper step to explore why training on longer CoTs leads to a decline in model’s reasoning performance. We randomly selected 100 problems from the tag model’s training set along with their responses under three different reasoning efforts. We first follow the existing study [3] to use gpt-4o to determine the total number of reasoning rounds contained in each response. Each reasoning round is defined as a complete reasoning process or verification process that contains a final answer. Besides, we further use gpt-4o to determine the number of reasoning rounds that contain erroneous steps or wrong final answers. The utilized prompt is shown in Appendix D. We visualize the average number of reasoning rounds, the average number of erroneous reasoning rounds, and the average ratio of erroneous reasoning rounds, on each problem under each reasoning effort in Figure 5. Note that all evaluated responses have correct final answers.
@@
Figure 5: Empirical results of loss masking on erroneous steps. Evaluation temperature is $0.$
First, we can see that the number of reasoning rounds consistently increases from low reasoning effort to high reasoning effort. It could lead to the overthinking issue in reasoning models [3], where redundant tokens are used to answer the question through repetitive verification. This, however, should not affect the accuracy if all reasoning rounds are correct. Unfortunately, we observe a pattern that the number and proportion of erroneous reasoning rounds also increase when the reasoning effort becomes higher. Training the model on more wrong steps would bring adverse effects to the model’s reasoning abilities, which can explain why scaling with high reasoning effort leads to worse results. Therefore, we can conclude that while including a certain incorrect and reflective steps can help the model learn to correct errors during inference, an excess of erroneous steps can have a detrimental impact on model’s learning.
To further validate the above claim, we conduct controlled validation experiments about training the model on erroneous reasoning steps with subsequent error corrections. Specifically, we first construct long CoTs with reflection and correction from scratch. We obtain initial CoTs with incorrect answers generated by LLaMA3.1-8B-Instruct on MATH training prompts, and use Qwen2.5-72B-Instruct to critique them (i.e., identify the erroneous step and generate suggestions for correction). Then, we enable LLaMA3.1-8B-Instruct to generate self-reflections and corrections based on the critiques, and merge these with the initial responses to form long CoTs. We fine-tune LLaMA3.1-8B-Instruct on the created long CoTs. In the results shown in Figure 5, we can observe that if we apply loss masking to the tokens in the identified wrong steps (i.e., not calculating loss on these erroneous steps), the performance is better compared to calculating loss on all steps/tokens. This further validates our claim that training on erroneous steps can negatively impact model performance. Additional experiments and empirical analysis are in Appendix I.